7. What is the energy released in this alpha decay reaction 83 Bi 20ST1+ He? he atomic mass of 212 is 211.9912857 u and that of 208 is 212 (T Bi 207.9820187 u) MeV
Answer #1

Q = mfc2 - mic2

Q = (211.9912857 u - 207.9820187 u- 4.002602u )c2

Q =( 6.665 * 10-3 u)c2

Im MeV -

Q = ( 6.665 * 10-3 u)c2   * (931.494 MeV/c2)/u)

Q = 6.02840 MeV

Thank you !!! Have fun